OpenStudy (anonymous):

integral of sin(x)^2/cosx

OpenStudy (anonymous):

you might notice that \(1-\sin^2 x = cos^2 x\), so change the sin x in terms of cosine, before integrating.

OpenStudy (anonymous):

the integral of sec x is \(\ln|sec x+ tan x|+C\)